40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 04:08:53 2015, Jenne, Update, LSC, New Locking Paradigm? 
    Reply  Wed Feb 11 22:13:44 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes 
       Reply  Thu Feb 12 01:43:09 2015, rana, Update, LSC, New Locking Paradigm - LSC model changes TF.gif
       Reply  Thu Feb 12 11:14:29 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra ALS_REFL_comboLockingCartoon_11Feb2015.PDFALS_REFL_comboLocking_11Feb2015.zip
          Reply  Thu Feb 12 11:59:58 2015, Koji, Update, LSC, New Locking Paradigm - Loop-gebra 
             Reply  Thu Feb 12 19:18:49 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra 
                Reply  Fri Feb 13 03:28:34 2015, rana, Update, LSC, New Locking Paradigm - Loop-gebra 
                   Reply  Sat Feb 14 20:20:24 2015, Jenne, Update, LSC, ALS fool cartoon ALSfool_LoopDiagram.pngALSfool_LoopDiagram.graffle.zip
                      Reply  Tue Feb 17 04:04:32 2015, Jenne, Update, LSC, ALS fool math OpenLoopGainComparison_16Feb2015.pngCancellationTFs_DecouplingIsArefl.png
                         Reply  Tue Feb 17 16:36:08 2015, Jenne, Update, LSC, ALS fool math ALSfool_LoopDiagram_17Feb2015.pngALS_REFL_comboLocking_16Feb2015.nb.zip
       Reply  Thu Feb 12 22:28:16 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes, screens modified 
          Reply  Sat Mar 7 19:15:17 2015, Jenne, Update, LSC, Modified zero crossing triggering LatchLogic_6March2015.png
    Reply  Thu Feb 12 03:43:54 2015, ericq, Update, LSC, 3F PRMI at zero ALS CARM Feb12_negativeTR.png
Message ID: 11009     Entry time: Thu Feb 12 01:43:09 2015     In reply to: 11007
Author: rana 
Type: Update 
Category: LSC 
Subject: New Locking Paradigm - LSC model changes 

With the Y Arm locked, we checked that we indeed can get loop decoupling using this technique.

The guess filter that we plugged in is a complex pole pair at 1 Hz. We guessed that the DC gain should be ~4.5 nm count. We then converted this number into Hz and then into deg(?) using some of Jenne's secret numbers. Then after measuring, we had to increase this number by 14.3 dB to an overall filter module gain of +9.3.

The RED trace is the usual 'open loop gain' measurement we make, but this time just on the LSC-MC path (which is the POY11_I -> ETMY path).

The BLUE trace is the TF between the ALS-Y phase tracker output and the FF cancellation signal. We want this to be equal ideally.

The GREEN trace is after the summing point of the ALS and the FF. So this would go to zero when the cancellation is perfect.

So, not bad for a first try. Looks like its good at DC and worse near the red loop UGF. It doesn't change much if I turn off the ALS loop (which I was running with ~10-15x lower than nominal gain just to keep it out of the picture). We need Jenne to think about the loop algebra a little more and give us our next filter shape iteration and then we should be good.

Attachment 1: TF.gif  5.632 MB  | Show | Show all
ELOG V3.1.3-